OpenStudy (anonymous):

Write the quadratic function in vertex form. y = x2 + 8x + 18

OpenStudy (anonymous):

i know you have to Subtract 18 from both sides: y - 18 = x^2 + 8x Complete the square by adding 16 to both sides: y - 18 + 16 = x^2 + 8x + 16 y - 2 = x^2 + 8x + 16 Express x^2 + 8x + 16 in binomial square form: y - 2 = (x + 4)^2 Add 2 to both sides: y = (x + 4)^2 + 2

OpenStudy (anonymous):

You can use the vertex formula to figure out the max height: t=−b2a t=−6402(−16) t=20 Next you'll need to insert t = 20 into the given equation: h(20) = - 16(20)^2 + 640(20)

OpenStudy (anonymous):

then simplify the expression on the right to solve for h(20)

OpenStudy (anonymous):

how would you solve for h(20)

OpenStudy (anonymous):

h(20) = - 16(20)^2 + 640(20)

OpenStudy (anonymous):

@saseal

OpenStudy (anonymous):

how did it turn from t into h?

OpenStudy (anonymous):

i have no idea im using this

OpenStudy (anonymous):

ok the general form of a parabola looks like this \[x^2=4py\] and the standard forms are the messy quadratic equations

OpenStudy (anonymous):

like that graph right?

OpenStudy (anonymous):

\[x^2+8x+18=0\]since you cant factorize this you need to complete the square\[x^2+8x=-18\]\[x^2+8x+(\frac{ 8 }{ 2 })^2=-18+(\frac{ 8 }{ 2 })^2\]\[x^2+8x+16=-18+16\]\[(x+4)^2=-2\]

OpenStudy (anonymous):

dont worry about that graph thing yet

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

@nono266

OpenStudy (anonymous):

you can understand it so far?

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

that means you dont understand...

OpenStudy (anonymous):

would it be - instead of the +

OpenStudy (anonymous):

\[(x+4)^2=-2\]\[(x+4)^2+2=0\]

OpenStudy (anonymous):

ugh yes!